# Here Be Dragons: Bimodal posteriors arise from numerical integration error in longitudinal models

### v1

@@ 2.1 Checking Canham
We use simulation to check whether the Canham model used in hmde is robust to posterior clustering due to numerical error with the RK45 method and adaptive step size. The differential equation defined in Canham et al., (2004) is
$f(Y(t),g_{max},y_{max},k)=g_{max}\exp\Bigg{(}-\frac{1}{2}\bigg{(}\frac{\ln(Y(t)/y_{max})}{k}\bigg{)}^{2}\Bigg{)},$ (8)
@@
which is extremely non-linear and has no analytic solution, necessitating numerical methods within the longitudinal model. Equation (8) has advantages for numerical estimation over the affine ODE: it is strictly positive which helps avoid the extreme oscillation we see in Figure 1 from negative growth, and it is bounded above by $g_{max}$ which limits how big positive growth increments can get. The affine function is unbounded above and below, and as a negative growth increment can lead to negative $Y(t)$ values even if the data is strictly positive (Canham is undefined for $Y(t)\leq 0$ due to the log function), it is possible for the numerical method to encounter extreme positive growth increments as well as negative ones.
There is no analytic solution to Equation (8) which would give access to ‘true’ sizes over time $Y(t)$. Instead we used the deSolve RK4 method with step size 0.0001 to project forward 50 units of time from an initial size of $Y(0)=1$. The true parameters $g_{max}=0.8,\ y_{max}=8,\ k=1$ were chosen as plausible values based on what was observed in O’Brien et al., (2024). To test whether the step size was small enough to give good numerical estimates, we did the same forward projection with step size 0.001 and compared the two, which gave a maximum difference of $3.02\times 10^{-13}$. We fit 10,000 chains to ‘observed’ sizes: from the smaller step size, we took 10 points, 5 years apart as the true values, then for each chain we add independent error from $\mathcal{N}(0,0.1)$, and rounded to precision 0.1. The scale of error and measurement precision aligns with real-world data from Condit et al., (2019). We did not use the more faithful error model of Rüger et al., (2011) which is a mixture of normal distributions because O’Brien et al., (2024) found that doing so does not meaningfully change estimation. We add error independently for each chain as we are interested in whether the Canham model has bias as well as multimodality because it is the function and numerical method implemented in hmde.
For analysis we look at the posterior distribution of parameter estimates. Scatter plots and histograms are used to examine clustering. We use the mean of the estimates and central 95% posterior credible interval to examine spread and bias. The posterior is unimodal and attempts to fit a mixture distribution with two clusters failed as a result, so we do not use one.
3 Results
We see persistent bimodality across numerical methods with fixed step size. The location of the second mode depends on the numerical method. Changing the location of the parameter priors is unable to constrain the occurrence of the erroneous mode, while constraining the prior spread only eliminates the second mode when the standard deviation is too small to be of practical use. The methods implemented in hmde produce unbiased, unimodal estimates for both the affine and Canham models.
3.1 Numerical methods have persistent bimodality
All the step sizes tested with the RK4 algorithm had a second cluster of erroneous parameter estimates. As the step size shrinks, the erroneous cluster moves further away from the true values, and the probability of erroneous convergence decreases as shown in Table 3.1. Figure 4 shows the cluster distance in Panel (b), where all 30,000 estimates are plotted together. There is a stack of points at (9.95, 0.991) which is the first cluster of estimates for all step sizes, and each step size has a second cluster spread further from that point as the step size decreases. The clusters all lie along a line preserving the asymptotic size at $\hat{\beta}_{0}/\hat{\beta}_{1}\approx 10$, which is also visible in Figure 4(c) and (d). Panel (c) shows Equation (1) with parameters fit to the mean of each erroneous cluster compared to the true values, and Panel (d) shows the analytic solution Equation (3) with the same. The erroneous cluster has much faster analytic solution convergence to the asymptotic size, but as we saw in Figure 1, that leads to bad numerical behaviour.
@@ 3.1 Numerical methods have persistent bimodality
Table 1: Posterior finite mixture model estimates for MCMC testing different numerical methods. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$.
@@
Numerical method Step size $\mu_{\beta_{0},1}$ $\mu_{\beta_{1},1}$ $\mu_{\beta_{0},2}$ $\mu_{\beta_{1},2}$ prob.
RK4 step size analysis 0.5 9.956 0.9917 49.26 4.913 0.3195
0.25 9.949 0.991 104.5 10.47 0.1044
0.125 9.949 0.9909 211.6 21.57 0.0254
@@
Figure 6 gives a close-up view of the two clusters for each step size, and show that there is variation in the estimates at the small scale that is not visible in Figure 4(b). There is a strong relationship between the parameters that reflects the asymptotic size being tightly constrained by the observations at $\hat{\beta}_{0}/\hat{\beta}_{1}\approx 10$.
Figure 4: (a) gives the simulated data with measurement error and precision of 0.1 compared to the analytic solution for $\beta_{0}=10$, $\beta_{1}=1$. Outside of the specified situations with independent error for each chain, estimation is based on the same data. (b) is a scatter plot of all posterior estimates, good and bad, for the default priors across three step sizes using RK4. The ‘points’ are distributions of estimates. There is a stacked cluster for all step sizes at the true values, and each step size has a separate second cluster of erroneous values. (c) gives the erroneous differential equations and shows that all estimates converge to the same maximum size. (d) shows the analytic solutions using the erroneous parameter combinations to show that the erroneous values give sizes that converge to the asymptotic maximum much faster than the true solution.
Figure 5: Estimated sizes over time across three step sizes compared to the true model for simulated data, and numerical error. (a), (c), and (e) show that the estimated sizes over time for the erroneous cluster align closely with the analytic solution for the true $\beta$s across the step sizes. The slight bias for step size 0.125 is likely due to the prior constraining $\hat{\beta}_{0}$. (b), (d), and (f) show the numerical error for each erroneous cluster, where the initial step has the largest error and then there is convergence to the same asymptotic size.
Figure 6: Posterior estimates scatter plots across three step sizes. (a) and (b) show clusters for step size 0.5. (c) and (d) show the clusters for step size 0.25, where the second cluster is further away from the first than for the larger step size and occurs less frequently. (e) and (f) show the first, and even more distant and less likely second cluster for the smallest step size 0.125.
We tested a better numerical method as well – RK45 with adaptive step size – and found that under some prior specifications bimodality still occurred. Figure 8 shows unimodal posteriors for both the analytic and RK45 posteriors when $\beta_{c}$ has a normal prior, while Figure 8 shows the bimodal posterior that arose when we were testing a log-normal distribution for $\beta_{c}$ with RK45, a problem that the analytic solution does not have. Log-normal priors are common so we include this example as a warning, even though the prior implemented in hmde does not demonstrate bimodality with RK45. Table 3.1 gives the mixture-model analysis for RK45. We do not consider the two clusters for $\beta_{c}\sim\log\mathcal{N}$ to be meaningfully different as they are so close together.
3.2 Changes to estimation methods do not address the underlying issue
Table 3.3 shows that all alternate priors except for one with an unreasonably small standard deviation picked up an erroneous second mode. The prior density is calculated after the fact by taking the product of normal distribution function values (as the two parameters are treated as independent in the prior) at the posterior cluster’s location using the prior parameters for mean and standard deviation. The density represents the height of the prior probability density function at that point. The posterior probability comes from the mixture distribution estimate.
Changing the prior central location had minimal impact, as can be seen in the first three prior configurations in Table 3.3. The ‘good’ modes had very similar locations, and the probability of converging to the second (erroneous) mode was not substantially different across the fits. The behaviour of the different centres is very similar the RK4 with 0.5 step size and default prior as seen in Table 3.1. That the posterior probability did not change even when the prior was centred at one or the other cluster indicates that changing the prior location does not help with constraining erroneous convergence.
A much larger prior standard deviation of 10 produced results that were not meaningfully different from the default of 2. Reducing the prior standard deviation to 0.1 changed where the second mode was, but did not meaningfully constrain its occurrence with a reduction from 32% to 26%. The even more unreasonably small prior standard deviation (0.001) was able to eliminate the second mode as the distance between the two means from the mixture model was vanishingly small. The purpose of including an unreasonable standard deviation prior was to demonstrate that it is mathematically possible to constrain where the posterior goes, but to do so requires something that would not be used in practice.
3.3 Our chosen methods produce unbiased, unimodal estimates
The simulation tests for the affine model with an analytic solution produced a unimodal distribution that was centred within 0.3% of the true values as shown in Table 3. Using the analytic solution means we do not introduce the non-identifiability that arises from numerical error so this result is in line with our theory.
Table 2: Posterior finite mixture model estimates for MCMC with different priors. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. The numerical integration was RK4 with a step size of 0.5. The $\beta_{c}$ prior mean includes the $\bar{y}$ shift, and the $\beta_{1}$ mean has a log transformation.
@@ 3.3 Our chosen methods produce unbiased, unimodal estimates
The overall behaviour of the posterior estimation for Canham is shown in Figure 10. We have a unimodal posterior distribution with no evidence of additional clusters in any dimension. Parameter estimation information is given in Table 3. We can be reasonably confident that the Canham function is not subject to the same pathologies as Equation (1) given the RK45 algorithm with adaptive step size used in the longitudinal model.
@@
Table 3: The affine model with an analytic solution and the Canham model with RK45 and an adaptive step size show no evidence of bias.
@@ 4 Discussion
This paper provides theoretical and empirical evidence of a bimodal posterior arising from numerical error in a longitudinal model for an affine first order ODE. One posterior mode aligns with good estimates of the true parameter values for simulated data, and at that mode the numerical method is reasonably well behaved. The other mode has extremely bad estimates of the true parameters, and badly behaved numerics as well. Theoretically, we show that in our chosen situation the RK4 numerical method has two parameter values that give the same sizes over time. In the empirical results, we see two clusters of parameter estimates that produce statistically indistinguishable $\hat{Y}(t_{j})$. The empirical parameter estimates align closely with the theoretical results.
We demonstrated the bimodal posterior for Equation (1), and show that the behaviour can occur for different step sizes, priors, true parameters, numerical methods, and estimation methods. Where the second mode is located depends on the numerical method and step size, as which combination of bad parameters gives ‘correct’ $\hat{Y}(t_{j})$ depends on both those factors. Changes to the prior distribution were unable to appropriately constrain the second mode, but implementing the analytic solution instead of using a numerical method eliminated the erroneous parameters in MCMC sampling.
The persistence of bimodal posteriors due to numerical error indicate that choice of integration method matters beyond bias in estimating the posterior parameters. Bias has been tested in Agapiou et al., (2014) and O’Brien et al., (2024). The integration methods chosen to be implemented in the package hmde produced unimodal, unbiased estimates when tested in simulation for both the affine model where we used the analytic solution, and the second non-linear DE known as the Canham function (Canham et al.,, 2004) which under RK45 with an adaptive step size has unbiased, unimodal posteriors.
@@
Estimation methods are a consideration for models for both material (computation time) and theoretical reasons. We tested the computationally-intensive MCMC sampling method, and a computationally inexpensive but demonstrably less robust L-BFGS optimisation method that maximises the joint posterior likelihood. Neither avoided bimodality on their own because they too do not address the underlying numerical problems. MCMC sampling requires the numerical calculations to be run thousands of times so while an erroneous cluster can be identified after the fact with a high precision (and computationally expensive) numerical method, implementing that directly in the longitudinal model is not so feasible. That problems can be identified after the fact is helpful at least. In some cases multiple chains run in parallel can detect a bimodal posterior if they converge to different points, which is how we first observed the issue as tweaks to the von Bertalanffy model in hmde with RK4 failed to avoid the occasional chain converging to a secondary mode. Working with a small number of parallel chains is less reliable as a means of identifying posterior bimodality than large-scale simulation testing as we have done in this paper, particularly when the probability of converging to the second mode is small as for the affine model with step size 0.125.
@@
The biggest consequence of bimodal non-identifiability in Bayesian inverse methods that use numerical integration is that even if a reasonable set of ‘true’ parameters is known, stability for the numerical methods matters for potentially a very large region of the parameter space. For the smallest step size we tried, the Euclidean distance between the good and bad modes was 217, and we found that constraining the priors did not meaningfully affect the posterior mode probability for the larger step size.
The numerical pathologies we found are adjacent to those addressed in the typical Bayesian workflow (Gelman et al.,, 2020), as the error comes from the longitudinal structure that is not part of most models. The simulation testing that we propose for posterior bimodality is relatively easy to implement, compared to hard-core numerical analysis for example, and can be done prior to large-scale dataset estimation. Aside from the affine model, we applied simulation tests to the extremely non-linear Equation (8), where we cannot apply an analytic solution to avoid bimodality as there isn’t one. In the case of the Canham function we are able to test for evidence of a bimodal posterior and parameter bias under the implemented numerical method – RK45 with adaptive step size – and find none. As a result we can be confident that the implementation of the Canham model in hmde is unlikely to be as problematic as the von Bertalanffy model was without the analytic solution.
@@
Non-identifiability arising from numerical error in a longitudinal model does not appear to be addressed elsewhere in the identifiability literature. More attention is paid to situations where the data is insufficient to uniquely specify estimates from a theoretical standpoint, whether due to too large a parameter space for the observations (Gelfand and Sahu,, 1999) or large measurement uncertainty (Auger-Méthé et al.,, 2016). Gelman et al., (2020) Section 11 addresses multi-modality that arises from orbit periodicity where there are not enough observations to select a specific value. Our problem arises in the model, but the literature around model structure tends to focus on over-parameterisation (Rannala,, 2002), or prior specification (Gelfand and Sahu,, 1999), neither of which are the problem here. As such, we hope this case study can expand the known issues with identifiability in inverse problems and Bayesian methods, and provide answers for someone else who has gotten stuck on a similar issue.

### v2

@@ 2.3.1 Checking Canham
We use simulation to check whether the Canham model is robust to posterior clustering due to numerical error with the RK45 method and adaptive step size. The differential equation defined in Canham et al., (2004) is
$f(Y(t),g_{max},y_{max},k)=g_{max}\exp\Bigg{(}-\frac{1}{2}\bigg{(}\frac{\ln(Y(t)/y_{max})}{k}\bigg{)}^{2}\Bigg{)},$ (2.20)
@@
which is extremely non-linear and has no analytic solution, necessitating numerical methods within the longitudinal model. Equation (2.20) has advantages for numerical estimation over the affine ODE: it is strictly positive which helps avoid the extreme oscillation we see in Figure 1 from negative growth, and it is bounded above by $g_{max}$ which limits how big positive growth increments can get. The affine function is unbounded above and below, and because a negative growth increment can lead to negative $Y(t)$ values even if the data is strictly positive (Canham is undefined for $Y(t)\leq 0$ due to the log function), it is possible for the numerical method to encounter extreme positive growth increments as well as negative ones.
@@
Table 1: Posterior finite mixture model estimates for MCMC testing different numerical methods for the affine ODE. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. The sample standard deviations of each cluster for each parameter are given in parentheses. Note that for each choice of step size, RK4 converged to one of two very different solutions.
@@
Numerical method Step size $\mu_{\beta_{0},1}\ (sd(\hat{\beta}_{0},1))$ $\mu_{\beta_{1},1}\ (sd(\hat{\beta}_{1},1))$ $\mu_{\beta_{0},2}\ (sd(\hat{\beta}_{0},2))$ $\mu_{\beta_{1},2}\ (sd(\hat{\beta}_{1},2))$ prob.
RK4 step size analysis 0.5 9.956 (0.011) 0.9917 (0.0012) 49.26 (0.015) 4.913 (0.00075) 0.3195
0.25 9.949 (0.011) 0.991 (0.0012) 104.5 (0.024) 10.47 (0.00072) 0.1044
0.125 9.949 (0.011) 0.9909 (0.0011) 211.6 (0.038) 21.57 (0.00067) 0.0254
@@
There is no analytic solution to Equation (2.20) which would give access to ‘true’ sizes over time $Y(t)$. Instead we used the deSolve RK4 method with step size 0.0001 to project forward 50 units of time from an initial size of $Y(0)=1$. The true parameters $g_{max}=0.8,\ y_{max}=8,\ k=1$ were chosen as plausible values based on what was observed in O’Brien et al., (2024). To test whether the step size was small enough to give good numerical estimates, we did the same forward projection with step size 0.001 and compared the two, which gave a maximum difference of $3.02\times 10^{-13}$. We fit 10,000 chains to ‘observed’ sizes: from the smaller step size, we took 10 points, 5 years apart as the true values, then for each chain we add independent error from $\mathcal{N}(0,0.1)$, and rounded to precision 0.1. We added error independently for each chain as We are interested in whether the Canham model has evidence of bias as well as multi-modality.
We use visualisations of the distribution of posterior means. Scatter plots and histograms are used to examine clustering. We use the mean of the estimates and central 95% posterior credible interval to examine spread and bias. The posterior is unimodal and attempts to fit a mixture distribution with two clusters failed as a result, so we do not use one.
3 Affine DE results
We see persistent bimodality across numerical methods with fixed step size. The location of the second mode depends on the numerical method. Changing the location of the parameter priors does not prevent convergence to the erroneous mode, while constraining the prior spread only eliminates the second mode when the standard deviation is too small to be of practical use. The analytic solution produces accurate, unimodal posteriors.
3.1 Numerical methods have persistent bimodality
All the step sizes tested with the RK4 algorithm had a second cluster of erroneous parameter estimates. As the step size shrinks, the erroneous cluster moves further away from the true values, and the probability of erroneous convergence decreases as shown in Table 2.3.1. Figure 3 shows the cluster distance in Panel (b), where all 30,000 estimates are plotted together yet they appear as if single points due to the tiny within-cluster variance, as in Table 2.3.1. There is a stack of points at (9.95, 0.991) which is the first cluster of estimates for all step sizes, and each step size has a second cluster spread further from that point as the step size decreases. The clusters all lie along a line preserving the asymptotic size at $\hat{\beta}_{0}/\hat{\beta}_{1}\approx 10$, which is also visible in Figure 3(c) and (d). Panel (c) shows Equation (2.1) with parameters fit to the mean of each erroneous cluster compared to the true values, and Panel (d) shows the analytic solution Equation (2.3) with the same. The erroneous cluster has much faster analytic solution convergence to the asymptotic size, but as we saw in Figure 1, that leads to bad numerical behaviour.
Table 2: Posterior finite mixture model estimates and prior densities for MCMC with different priors. The prior density is the density of the prior distribution at the cluster centre and indicates relative probability before model fitting, while posterior probability is based on the number of points in each cluster. True parameter values are $\beta_{0}=10$ and $\beta_{1}=1$. Numerical integration used the RK4 method with a step size of 0.5. The $\beta_{c}$ prior mean includes the $\bar{y}$ shift, and the $\beta_{1}$ mean has a log transformation. Note that results, where the posterior mean was one of two very different solutions (each with high probability), are robust to prior choice, with the second solution only disappearing when an extremely tight prior was chosen around the first solution.
@@ 3.1 Numerical methods have persistent bimodality
Figure 5: Estimated $\hat{Y}(t)$ across three step sizes compared to the true model for simulated data, and numerical error. (a), (c), and (e) show that the estimated sizes over time for the erroneous cluster align closely with the analytic solution for the true $\beta$s across the step sizes. The slight difference in asymptotic size for step size 0.125 is likely due to the prior constraining $\hat{\beta}_{0}$. (b), (d), and (f) show the numerical error for each erroneous cluster, where the initial step has the largest error and then there is convergence to the same asymptotic size.
We tested a better numerical method as well – RK45 with adaptive step size – and found that under some prior specifications bimodality still occurred. Figure 6 shows unimodal posteriors for both the analytic and RK45 posteriors when $\beta_{c}$ has a normal prior, while Figure 7 shows the bimodal posterior that arose when we was testing a log-normal distribution for $\beta_{c}$ with RK45, a problem that the analytic solution does not have. Log-normal priors are common so we include this example as a warning. Table 2.3.1 gives the mixture-model analysis for RK45. We do not consider the two clusters for $\beta_{c}\sim\log\mathcal{N}$ to be meaningfully different as they are so close together.
3.2 Changes to estimation methods do not address the underlying issue
Table 3.1 shows that all alternate priors except for one with an unreasonably small standard deviation picked up an erroneous second mode. Changing the prior central location had minimal impact, as can be seen in the first three prior configurations in Table 3.1. The ‘good’ modes had very similar locations, and the probability of converging to the second (erroneous) mode was not substantially different across the fits. The behaviour of the different centred priors is very similar to the RK4 with 0.5 step size and default prior as seen in Table 2.3.1. That the posterior probability did not change even when the prior was centred at one or the other cluster indicates that changing the prior location does not help with constraining erroneous convergence.
A larger prior standard deviation of 10 produced results that were not meaningfully different from the default of 2. Reducing the prior standard deviation to 0.1 changed where the second mode was, but did not meaningfully constrain its occurrence with a reduction from 32% to 26%. The even more unreasonably small prior standard deviation (0.001) was able to eliminate the second mode as the distance between the two means from the mixture model was vanishingly small. The purpose of including an unreasonable standard deviation prior was to demonstrate that it is mathematically possible to constrain where the posterior goes, but to do so requires something that would not be used in practice.
3.3 Our chosen methods produce accurate, unimodal estimates
The simulation tests for the affine model with an analytic solution produced a unimodal distribution that was centred within 0.3% of the true values as shown in Table 3 which is a negligible empirical bias. Using the analytic solution means we do not introduce the non-identifiability that arises from numerical error so this result is in line with our theory.
4 Canham results
The overall behaviour of the posterior estimation for Canham is shown in Figure 9. We have a unimodal posterior distribution with no evidence of additional clusters in any dimension. Parameter estimation information is given in Table 3. We can be reasonably confident that the Canham function is not subject to the same pathologies as Equation (2.1) given the RK45 algorithm with adaptive step size used in the longitudinal model.
@@ 4 Canham results
Table 3: The affine model with an analytic solution and the Canham model with RK45 and an adaptive step size show minimal evidence of bias.
@@ 5 Discussion
This paper provides theoretical and empirical evidence of a multi-modal posterior arising from numerical error in a longitudinal model for an affine first order ODE. One posterior mode aligns with good estimates of the true parameter values for simulated data, and at that mode the numerical method is reasonably well behaved. The other mode has extremely bad estimates of the true parameters, and badly behaved numerics as well. The existence of non-identifiability for RK4 in the case of an affine ODE demonstrates that care must be taken when using numerical methods for longitudinal models.
We show that the affine model can use an analytic solution to avoid the numerical pathology, and also that in the case of the Canham model where there is no analytic solution to use, the choice of RK45 with adaptive step size is sufficient as we have no evidence of multi-modal posteriors. The simulation testing procedure is more generally applicable and should form part of longitudinal model implementation if there is evidence of bad posterior behaviour such as divergent chains. It is also possible to test estimates from real world data after the fact by using a more accurate numerical method than the one implemented in the model.
Using numerical theory, we proved that in general, the RK4 algorithm in a longitudinal model for affine first-order ODEs with negative slope will produce multi-modal results, and demonstrate with a specific ODE. In the empirical results, we see two clusters of parameter estimates that produce statistically indistinguishable $\hat{Y}(t_{j})$. The empirical parameter estimates align closely with the theoretical results.
For the end user, the process of testing common ways to adapt a model and prevent problems will hopefully be demonstrative. We tested different step sizes for the numerical method, different priors, estimation processes, and implementations of the longitudinal structure. Identifying that the divergent chains problem originated from the numerical method was difficult, but the symptom was distinct once recognised: independent chains converging to different, but consistent, parameter values, one of which was a good estimate of the true parameters. If this behaviour is observed in a longitudinal model using numerical methods, check how the numerics behave at the different parameter values. Test both the numerical method as implemented in the longitudinal model, and a (computationally more expensive) more accurate method as well.
We demonstrated the bimodal posterior for Equation (2.1), and show that the behaviour can occur for different step sizes, priors, true parameters, numerical methods, and estimation methods. Where the second mode is located depends on the numerical method and step size, as which combination of bad parameters gives ‘correct’ $\hat{Y}(t_{j})$ depends on both those factors. Changes to the prior distribution were unable to appropriately constrain the second mode, but implementing the analytic solution instead of using a numerical method eliminated the erroneous parameters in MCMC sampling.
The persistence of bimodal posteriors due to numerical error indicate that choice of integration method matters beyond bias in estimating the posterior parameters. Bias has been tested in Agapiou et al., (2014) and O’Brien et al., (2024). We are able to select integration methods that produce unimodal estimates with negligible bias. For the affine model we used the analytic solution, and for the second non-linear DE known as the Canham function (Canham et al.,, 2004) RK45 with an adaptive step size produces accurate, unimodal posteriors.
@@
Estimation methods are a consideration for models for both material (computation time) and theoretical reasons. We tested the computationally-intensive MCMC sampling method, and a computationally inexpensive but demonstrably less robust L-BFGS optimisation method that maximises the joint posterior likelihood. Neither avoided bimodality on their own because they too do not address the underlying numerical problems. MCMC sampling requires the numerical calculations to be run thousands of times so while an erroneous cluster can be identified after the fact with a high precision (and computationally expensive) numerical method, implementing that directly in the longitudinal model is not so feasible. That problems can be identified after the fact is helpful at least. In some cases multiple chains run in parallel can detect a bimodal posterior if they converge to different points, which is how we first observed the issue as tweaks to the von Bertalanffy model with RK4 failed to avoid the occasional chain converging to a secondary mode. Working with a small number of parallel chains is less reliable as a means of identifying posterior bimodality than large-scale simulation testing as we have done in this paper, particularly when the probability of converging to the second mode is small as for the affine model with step size 0.125.
@@
The biggest consequence of this form of non-identifiability in Bayesian inverse methods that use numerical integration is that even if a reasonable set of ‘true’ parameters is known, stability for the numerical methods matters for potentially a very large region of the parameter space. For the smallest step size we tried, the Euclidean distance between the good and bad modes was 217, and we found that constraining the priors did not meaningfully affect the posterior mode probability for the larger step size.
The numerical pathologies we found are adjacent to those addressed in the typical Bayesian workflow (Gelman et al.,, 2020), as the error comes from the longitudinal structure that is not part of most models. The simulation testing that we propose for posterior bimodality is relatively easy to implement, compared to hard-core numerical analysis for example, and can be done prior to large-scale dataset estimation. Aside from the affine model, we applied simulation tests to the extremely non-linear Equation (2.20), where we cannot apply an analytic solution to avoid bimodality as there isn’t one. In the case of the Canham function We are able to test for evidence of a bimodal posterior and parameter bias under the implemented numerical method – RK45 with adaptive step size – and find none. As a result we can be confident that the implementation of the Canham model using RK45 is unlikely to be as problematic as the von Bertalanffy model was without the analytic solution.
@@
Non-identifiability arising from numerical error in a longitudinal model does not appear to be addressed elsewhere in the identifiability literature. More attention is paid to situations where the data is insufficient to uniquely specify estimates from a theoretical standpoint, whether due to too large a parameter space for the observations (Gelfand and Sahu,, 1999) or large measurement uncertainty (Auger-Méthé et al.,, 2016). Gelman et al., (2020) Section 11 addresses multi-modality that arises from orbit periodicity where there are not enough observations to select a specific value. Our problem arises in the longitudinal part of the model structure, but the literature around model structure tends to focus on over-parameterisation (Rannala,, 2002), or prior specification (Gelfand and Sahu,, 1999), neither of which are the problem here. As such, we hope this case study can expand the known issues with identifiability in inverse problems and Bayesian methods, and provide answers for someone else who has gotten stuck on a similar issue.